te ti ti 7 presents a method for modeling this problem. u 5 LER, Kam . offers useful background for problems like this one. A cYlin er is rotating about an axis that passes through the center of each circular end piece. The cylinder has a radius of 0.0650 m, an angular speed ot 6 .0 ad/s, and a mo ment o ne tia ot 0 65 ka- 2. A brake shoe presses against the su tace ot the c nder and applies a tangential trictional torce to it. The trictional torce educes he angular s eed or the cylinder by a tactor o 7 during a time of 0.00 s. (a) Find the magnitude of the angular deceleration ot the cylinder. (b) Find the magnitude of the force of triction appied by the brake shoe. of0.005, (a) Find the meritude otradananlardeomer torsnertilem0.565kpCwusbrExauul,AQotters us nr dratien of the cvinde. ful background for pr hetat radys 2 (b) Numbar
Answer #1

a] a = [v2-v1]/t = [88 - 88/7]/8 = 9.53 rad/s^2

b] F = torque/R = [0.565*9.53]/0.065 = 82.8 N